### v1

@@ 5 Conclusions
The dataset is private, but comprehensive information is provided in Sutter et al. (2025b). The code is available at https://github.com/cgsutter/DRIVE-clean.
@@ Acknowledgments
This material is based upon work supported by the U.S. National Science Foundation under Grant No. RISE-2019758. This work is part of the NSF AI Institute for Research on Trustworthy AI in Weather, Climate, and Coastal Oceanography (NSF AI2ES). This research is also made possible by the New York State (NYS) Mesonet. Original funding for the NYS Mesonet (NYSM) buildup was provided by Federal Emergency Management Agency grant FEMA-4085-DR-NY. The continued operation and maintenance of the NYSM is supported by National Mesonet Program, University at Albany, Federal and private grants, and others. The xCITE lab (XCITE, 2025) provided computational resources for model development.

### v2

@@ 5 Conclusion and future work
The dataset is private, but comprehensive information is provided in Sutterdatasheet. The code is available at https://github.com/cgsutter/DRIVE-clean.
@@ Acknowledgments
This material is based upon work supported by the U.S. National Science Foundation under Grant No. RISE-2019758. This work is part of the NSF AI Institute for Research on Trustworthy AI in Weather, Climate, and Coastal Oceanography (NSF AI2ES). xcite provided computational resources for model development. This research is also made possible by the New York State (NYS) Mesonet. Original funding for the NYS Mesonet (NYSM) buildup was provided by Federal Emergency Management Agency grant FEMA-4085-DR-NY. The continued operation and maintenance of the NYSM is supported by National Mesonet Program, University at Albany, Federal and private grants, and others.